OpenStudy (anonymous):

whats the answer to 12x+6x+9x-3y+-7y+y

OpenStudy (dmezzullo):

Wouldn't u just add all the like term?

OpenStudy (dmezzullo):

terms**

OpenStudy (anonymous):

I did but I need someone to teach me step by step of doing the process

OpenStudy (dmezzullo):

Ok so u can only add terms that are the same as in 12x+1x+12yx U can only add the 12x+1x and the other stays alone u know that right?

OpenStudy (anonymous):

I Know How How To Add The Varibels But Why Is There A y And X??

OpenStudy (dmezzullo):

Bc they are variables. If they just put 1+2+3 that wouldn't put any challenge in the problem and you'd learn nothing

OpenStudy (anonymous):

This Is The Answer I Got But I Don't Know If Its Correct I Got 27x + -5y

OpenStudy (dmezzullo):

Lemme check

OpenStudy (anonymous):

Ok

OpenStudy (dmezzullo):

Yea that is correct :3

OpenStudy (anonymous):

Really? :o

OpenStudy (dmezzullo):

Yes!

OpenStudy (dmezzullo):

Good Job :)))

OpenStudy (anonymous):

Ok I Need Help On This Othere One

OpenStudy (dmezzullo):

Ok

OpenStudy (anonymous):

3(-4+7y) -3x(2+3y) Ok I Got -4x+23y Is It Correct Or Not?

OpenStudy (dmezzullo):

I wanna say its not right but ima get a second opion ok?

OpenStudy (dmezzullo):

@LoveYou*69

OpenStudy (anonymous):

Alright Then

OpenStudy (anonymous):

3(-4+7y) -3x(2+3y) -12 + 21y - 6x - 9xy <--thats what I got

OpenStudy (anonymous):

But How?

OpenStudy (anonymous):

Did You Add The Varibels?

OpenStudy (anonymous):

Foil method, 3(-4 + 7y) -3x(2 + 3y) Lets do this part first: 3(-4 + 7y) 3•-4 = -12 3•7y = 21y Now: -3x(2 + 3y) -3x•2 = -6x -3x•3y = -9xy So you have -12 + 21y - 6x - 9xy

OpenStudy (anonymous):

Oh Ok I Get It Now Thanks Cx

OpenStudy (anonymous):

Anytime(:

OpenStudy (anonymous):

12x+6x+9x-3y+-7y+y Combine all x's: 2x+6x+9x= ? Combine all the y's: -3y+y-7= ? Take those answers and put them together. 17x-9y Is what I got, because 2x+6x+9x=17x and -3y+y-7=-9y Does this help? \(\Huge\color{darkorchid}{\ddot\smile}\)

OpenStudy (anonymous):

Could You Help Me With This Other Problem Love You *69?

OpenStudy (anonymous):

Sure, I'll try

OpenStudy (anonymous):

It's 27, not 17. My bad.

OpenStudy (anonymous):

Ok 4(x+5y)+(5x+7) I Got 9x+10y

OpenStudy (anonymous):

Is It Correct or Wrong?

OpenStudy (anonymous):

4(x+5y)+(5x+7) 4x + 20y + 5x + 7 9x + 20y + 7 <--thats what I got

OpenStudy (anonymous):

Ok so I Got The 9x Right But Not The Rest

OpenStudy (anonymous):

Yep, you didn't get 4•5y = 20y and the 7

OpenStudy (anonymous):

Oh Ok Then Thanks Could You Help Me With This Other One?

OpenStudy (anonymous):

lol, sure

OpenStudy (anonymous):

Cynthia, close this question and open a new one. We don't want to spam this post. c:

OpenStudy (anonymous):

Oh But I Don't Know How I Barley Started This?

OpenStudy (anonymous):

Right under your question there should be two buttons, one says bump and the other says close. Click close. c:

OpenStudy (anonymous):

Loveyou *69 Did You Get The Message?

OpenStudy (anonymous):

\(\Huge\bf\color{darkorchid}{Welcome\ To\ OpenStudy!\ :D}\)